OpenStudy (anonymous):

how do you graph y = x – 2 using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (mosaic):

In y = mx + b, m is the slope and b is the y-intercept. Compare this to y = x - 2. Can you tell me what is the slope and what is the y-intercept?

OpenStudy (mathmale):

If you saw y = x - 2 on a quiz or test, what would you do if asked to graph this? Can you discern the slope of this straight line from y = x -2? The y-intercept? Think: y = mx + b

OpenStudy (anonymous):

y = x – 2 –2x + 3y = -6